# Projected Subgradient Ascent for Convex Maximization

### v1

@@ Abstract.
We consider the problem of maximizing a convex function over a closed convex set. Classical methods solve such problems using iterative schemes that repeatedly improve a solution. For linear maximization, we show that a single orthogonal projection suffices to obtain an approximate solution. For general convex functions over convex sets, we show that projected subgradient ascent converges to a first-order stationary point when using arbitrarily large step sizes. Taking the step size to infinity leads to the conditional gradient algorithm, and iterated linear optimization as a special case. We illustrate numerical experiments using a single projection for linear optimization in the elliptope, reducing the problem to the computation of a nearest correlation matrix.
@@ 1. Introduction
We consider the problem of maximizing a convex function over a nonempty closed and convex set $S$. Of particular interest is the case of linear optimization:
$\max_{x\in S}\;\langle c,x\rangle$ (1)
@@
Let $P_{S}(x)$ denote the unique point in $S$ closest to $x$. We show that $P_{S}(x_{0}+\eta c)$ converges to the unique optimal solution of (1) closest to $x_{0}$ as $\eta\to\infty$ (Theorem 2.7). We also give tight bounds on the quality of the solution obtained using finite $\eta$ (Lemma 2.4).
Beyond linear objectives, we investigate the use of projected subgradient ascent for maximizing convex functions. In the convex minimization setting, projected subgradient descent requires vanishing step sizes. In contrast, we show that for convex maximization, projected subgradient ascent converges to a first-order stationary point when using arbitrarily large step sizes (Theorem 3.14). This result holds in infinite-dimensional spaces, and it does not require differentiability, Lipschitz continuity of the gradient, or the Kurdyka-Łojasiewicz property (see, e.g. [2]).
In practice, the large step size regime of projected subgradient ascent may lead to faster convergence and yield meaningful behavior. In the limit, projected subgradient ascent is equivalent to the conditional gradient method with unit step size (Section 3.3). This limit also generalizes the iterated linear optimization paradigm introduced in [10].
The idea of using a single projection for linear optimization was previously considered in [16] in the context of linear programming. It was also considered recently, and independently of our work, in [24]. Compared to [24], the bounds shown here are tighter, and we prove convergence to a unique optimum solution besides proving convergence of the objective value.
@@
[7] undertook a complexity analysis and show that on many domains—the simplex, $\ell_{p}$-balls, nuclear-norm ball, the flow polytope, Birkhoff polytope, and permutahedron—the best known algorithms for linear optimization are asymptotically faster than the best known algorithms for orthogonal projection. The reduction of linear optimization to orthogonal projection gives further evidence that linear optimization over a convex set is never harder than projecting to the same set. While [7] and [24] suggest this as a negative result for projection-based methods, we emphasize that the reduction can be used to obtain positive results, in that efficient algorithms for projections lead to efficient algorithms for linear optimization.
In Section 4 we illustrate numerical experiments using projection for linear optimization in the elliptope, and the Goemans and Williamson [12] Max-Cut algorithm. We reduce the semidefinite programming problem to orthogonal projection and use an existing method to compute the nearest correlation matrix [22]. Our experiments show the projection method has comparable accuracy but is faster in practice when compared to SCS [20].
@@ 2. Single Projection for Linear Optimization
Since $S$ is closed and convex, and $\langle c,x\rangle$ is a linear function of $x$, the set of maximizers $\mathcal{M}(c)$ is non-empty, closed, and convex. Let $x_{0}\in\mathcal{H}$. Denote by $\|\cdot\|$ the norm induced by the inner product. By the Hilbert Projection Theorem, there is a unique solution in $\mathcal{M}(c)$ closest to $x_{0}$, which we denote
@@
We show that $x^{\eta}\to x^{*}$ as $\eta\to\infty$ and give explicit bounds on $\eta$ that guarantee a good approximation. This provides a method for approximating $x^{*}$ using a single orthogonal projection.
@@
The use of a single projection for linear optimization was considered in [16] in the context of linear programming, where projecting a suitably rescaled cost vector yields an optimal basic feasible solution. Mangasarian sought an explicit $\eta$ that produces an exact solution in the polyhedral setting. Here we consider more general (non-polyhedral) closed convex sets in a Hilbert space and derive bounds on $\eta$ that ensure an arbitrarily close approximation to the objective value. Note that for smooth sets, there is no finite $\eta$ that achieves optimality.
A result similar to Lemma 2.4 appears in [24]. However, our lemma provides a more precise characterization that depends on $x_{0}$ and we prove convergence of $P_{S}(x_{0}+\eta c)$ to a particular optimum solution, including in the case of general (non-polyhedral) convex sets $S$.
@@
The next key lemma shows $\langle c,x^{\eta}\rangle$ approaches $\langle c,x^{*}\rangle$ from below as $\eta$ grows.
@@
Proof.
Let $c=(\tfrac{1}{2\eta},1)$. The optimal solution is $x^{*}=(1,1)$ while $x^{\eta}=(\tfrac{1}{2},1)$. ∎
Figure 2. Illustration of Example 2.6. By selecting $\eta$ sufficiently large we can ensure $\langle c,x^{*}\rangle-\langle c,x^{\eta}\rangle\leq\epsilon$ independent of $c$, while $\|x^{*}-x^{\eta}\|$ remains large for some $c$.
@@
for all $\eta>0$. Therefore, $x^{\eta}$ lies in a closed ball of radius $\|x^{*}\|$ centered at $x_{0}$. Closed balls in Hilbert space are weakly sequentially compact. Consider an arbitrary increasing sequence $\{\eta_{j}\}$ such that $\{x^{\eta_{j}}\}\rightharpoonup\overline{x}$ for some $\overline{x}\in S$. We show that $x^{\eta_{j}}\to x^{*}$.
@@
Lemma 2.4 implies $\|x^{\eta}-x_{0}\|^{2}\leq\|x^{*}-x_{0}\|^{2}$. Therefore $\limsup_{j}\|x^{\eta_{j}}-x_{0}\|^{2}\leq\|x^{*}-x_{0}\|^{2}$. The weak lower semicontinuity of the norm yields $\|x^{*}-x_{0}\|\leq\liminf_{j}\|x^{\eta_{j}}-x_{0}\|$. Since the limit inferior is always at most the limit superior, combining both inequalities, we obtain
@@ 3. Projected Subgradient Ascent
In this section, we analyze the behavior of projected subgradient ascent for maximizing a convex function $f$ over a closed convex set $S\subseteq\mathcal{H}$. The problem of maximizing a convex function is NP‐hard (e.g., [21]). Therefore, we focus on obtaining first-order stationary points.
@@
converges to a global minimizer (see, e.g., [5]). In contrast, when maximizing a convex function using projected subgradient ascent (PGA) global convergence is no longer guaranteed, regardless of the choice of step sizes. Rather than vanishing steps, we focus on the case of arbitrarily large step sizes and show that PGA always converges to a first-order stationary point of $f$.
@@
Proposition 3.9 shows that in the convex maximization regime with PGA, the sequence $\{f(x_{k})\}$ is nondecreasing for any choice of step sizes. When $f$ is bounded, this implies $\{f(x_{k})\}$ converges. If the step sizes are bounded from above, Theorem 3.13 shows that the accumulation set of $\{x_{k}\}$ is connected. Moreover, if the step sizes are bounded from below, Theorem 3.14 shows all of the accumulation points of $\{x_{k}\}$ are first-order stationary points of $f$.
These results illustrate a contrast between convex minimization with PGD and convex maximization with PGA. The convex minimization setting requires vanishing steps for convergence, while in the convex maximization setting, convergence is guaranteed with large steps. In Section 3.3, we consider the limiting behavior when all the step sizes go to infinity.
@@ 3.1. Linear Functions
In the general case of possibly infinite-dimensional spaces, weak convergence under arbitrary step sizes can be shown using notions from monotone operator theory (see, e.g., [4]). The main result in this case is Theorem 3.7.
@@
Let $\{\eta_{k}\}_{k\in\mathbb{N}}$ be a sequence such that $\sum_{k=0}^{\infty}\eta_{k}=\infty$. Let $\{x_{k}\}_{k\in\mathbb{N}}$ be a sequence of iterates generated by PGA with $f(x)=\langle c,x\rangle$. The sequence converges weakly to a point in $\mathcal{M}(c)$.
@@
We emphasize the difference between Theorem 2.7 and Theorem 3.7. The first theorem analyzes the behavior of a single projected gradient step $P_{S}(x_{0}+\eta c)$ as the step $\eta\to\infty$, proving strong convergence to a unique solution. In contrast, the second theorem establishes weak convergence after an infinite number of finite steps.
@@ 3.2. Convex Functions
Now, we consider the case of maximizing a convex function with PGA.
@@
Previous results that apply to the setting of maximizing convex functions require significantly stronger assumptions. For example, the convergence results in [2] apply in finite dimensional spaces when $f$ is differentiable with a Lipschitz continuous gradient over a nonempty closed set $S$, and satisfies the Kurdyka-Łojasiewicz (KL) property.
@@
$f_{k}(x_{k+1})\geq f_{k}(x_{k}).$ (2)
Note that $x_{k+1}=P_{S}(x_{k}+\eta_{k}g_{k})$ and $x_{k}=P_{S}(x_{k})$. By Proposition 2.8 with $\eta_{1}=0$ and $\eta_{2}=\eta_{k}$ we have $\langle g_{k},x_{k+1}\rangle\geq\langle g_{k},x_{k}\rangle$. This implies $\langle x_{k+1}-x_{k},g_{k}\rangle\geq 0$ and (2) follows from the definition of $f_{k}$.
@@
Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\to\mathbb{R}$ be continuous and convex, $S\subseteq\mathcal{H}$ be a compact and convex set, and $\{x_{k}\}_{\mathbb{N}}$ be a sequence of iterates generated by PGA when the step sizes $\{\eta_{k}\}_{\mathbb{N}}$ satisfy $\limsup_{k\to\infty}\eta_{k}<\infty$. Then the set of accumulation points is connected.
@@
Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\to\mathbb{R}$ be continuous and convex, $S\subseteq\mathcal{H}$ be a compact and convex set, and $\{x_{k}\}_{\mathbb{N}}$ be a sequence of iterates generated by PGA when the step sizes $\{\eta_{k}\}_{\mathbb{N}}$ satisfy $\limsup_{k\to\infty}\eta_{k}<\infty$ and $\liminf_{k\to\infty}\eta_{k}>0$. Then every accumulation point of $\{x_{k}\}_{\mathbb{N}}$ is a first-order stationary point of $f$.
@@ 3.3. Conditional Gradient and Iterated Linear Optimization
$x_{k+1}=\operatorname*{\arg\!\max}_{z\in S}\,\langle x_{k},z\rangle,$
@@
4. Semidefinite Programming & Max-Cut
The Max-Cut problem seeks to partition the vertices of a graph into two disjoint sets that maximize the total weight of the edges crossing the partition.
Goemans and Williamson (GW) [12] introduced an efficient $0.878$-approximation algorithm for the Max-Cut problem based on a semidefinite programming relaxation followed by a randomized rounding step. The relaxation involves optimization over a convex body known as the elliptope.
The elliptope is the set of correlation matrices, defined as the set of positive semidefinite (PSD) matrices with unit diagonal entries,
$\mathcal{L}_{n}=\{X\in\mathbb{R}^{n\times n}\mathrel{\mathop{\ordinarycolon}}X\succeq 0,\operatorname{diag}(X)=\mathbf{1}\}.$
The GW relaxation involves linear optimization over the elliptope,
$X^{*}\in\operatorname*{\arg\!\max}_{X\in\mathcal{L}_{n}}\langle M,X\rangle,$ (7)
where M=WM=-W and $W\in\mathbb{R}^{n\times n}$ is the weighted adjecenty matrix of $G$.
The GW algorithm uses a randomized rounding procedure to produce a cut from $X^{*}$ with expected value at least $0.878$ times the value of the maximum cut.
Using the results in Section 2, we can find an approximate solution to (7) by projecting $\eta M$ onto $\mathcal{L}_{n}$,
$X^{\eta}=P_{\mathcal{L}_{n}}(\eta M)=\operatorname*{\arg\!\min}_{X\in\mathcal{L}_{n}}\|X-\eta M\|_{F},$ (8)
where $\|\cdot\|_{F}$ denotes the Frobenius norm.
Projecting to the elliptope (8) is known as the Nearest Correlation Matrix (NCM) problem. [13] uses Dykstra’s algorithm to solve the NCM problem and [15] applies Anderson acceleration to Dykstra. Qi and Sun (QS) [22] described a semismooth Newton method, which offers global convergence with local quadratic convergence.
A direct application of Lemma 2.4 yields a worst-case bound on the approximation error when using $X^{\eta}$ to approximate $X^{*}$,
$\langle M,X^{*}\rangle-\langle M,X^{\eta}\rangle\leq\frac{\|X^{*}\|_{F}^{2}}{2\eta}\leq\frac{n^{2}}{2\eta}.$
The final inequality holds because the squared norm of an $n\times n$ correlation matrix is bounded by $n^{2}$.
While interior-point methods can solve (7), their high computational cost limits their scalability. This has led to the adoption of more first-order alternatives like the Splitting Conic Solver [20]. Our numerical experiments show that for the Max-Cut SDP, the QS algorithm for the NCM problem produces competitive results when compared to SCS.
The numerical experiments were done with Python on a Linux computer with an Intel i7-13700 CPU @ 5.2 Ghz and 64GB of RAM. We implemented the QS algorithm using numpy and used the SCS implementation available in cvxpy. We evaluate both methods using the Gset dataset, a standard Max-Cut benchmark, focusing on graphs with 800 vertices. We set all algorithm parameters to their default setting and used $\eta=4000$ for the projection method. We restricted both implementations to use a single core for a fair comparison.
Table 1 compares the results using three metrics:
  • SDP objective: The inner product of the relaxed solution $\langle M,X\rangle$ [^1] Neither the QS algorithm nor the SCS produces a feasible solution. The QS algorithm outputs a PSD matrix, which we map to a correlation matrix by normalizing its Gram vector representation. The SCS algorithm outputs a symmetric matrix, which we map to a correlation matrix via the shrinking map defined in [14]: $X\to I+\tfrac{1}{1-\lambda}(X-\operatorname{diag}(X))$ with $\lambda=\lambda_{\min}(X-\operatorname{diag}(X)+I)$..
  • Cut value: The average Max-Cut value obtained by rounding the relaxed solution 100 times.
  • Runtime: The wall-clock time to solve the relaxed problem.
Of the twenty Gset graphs with 800 vertices, we randomly select ten to show in Table 1. Note how the SDP objective and cut values obtained using the two different methods are essentially indistinguishable. The key difference is runtime. The projection method is one to two orders of magnitude faster than SCS. For dense graphs (e.g., G1-G10), SCS takes 150-220 seconds. In contrast, the projection method takes approximately 18 seconds. The speedup is more pronounced on sparser graphs (e.g., G11-G20), where SCS can take over 7000 seconds. This provides some evidence that the projection approach is a scalable alternative for solving the Max-Cut SDP relaxation.
Runtime and Accuracy
Gset Graph Runtime (s) SDP Objective Cut Value
Projection SCS Projection SCS Projection SCS
G2 18.98 151.74 10005.7 10005.7 11274 11273
G3 18.76 179.85 9985.2 9985.2 11283 11277
G5 18.77 151.54 10047.5 10047.4 11290 11289
G7 18.36 215.44 10257.0 10256.9 1670 1662
G9 16.42 173.96 10242.9 10242.8 1707 1709
G11 64.59 7094.78 2447.4 2448.5 500 501
G14 17.34 1792.28 3378.0 3378.3 2927 2928
G18 24.38 1461.29 4535.8 4536.0 849 851
G19 22.24 1902.40 4553.6 4554.0 768 767
G20 29.60 1221.30 4537.3 4537.6 793 797
Table 1. Runtime and accuracy comparison between using NCM and SCS to solve instances of Max-Cut defined by the Gset graphs. The projection method returns a solution with comparable objective value and cut value, but is 8-100× faster on all instances.

### v2

@@ Abstract.
We consider the problem of maximizing a convex function over a closed convex set in a real Hilbert space. For linear functions, we show that a single orthogonal projection suffices to obtain an approximate solution. For continuous convex functions over convex sets, we show that projected subgradient ascent converges to a first-order stationary point when using arbitrarily large step sizes. Taking the step sizes to infinity leads to a deterministic variant of the conditional gradient algorithm, and iterated linear optimization as a special case.
Keywords: Linear optimization, Orthogonal projection, Nonlinear programming, Subgradient methods.
MSC: 90C05, 90C22, 90C25
@@ 1. Introduction
We consider the problem of maximizing a convex function over a nonempty closed and convex set $S$ in a real Hilbert space $\mathcal{H}$. Of particular interest is the case of linear optimization:
(1) $\max_{x\in S}\;\langle c,x\rangle$
@@
Let $P_{S}(x)$ denote the unique point in $S$ closest to $x$, i.e., the orthogonal projection. We show that $P_{S}(x_{0}+\eta c)$ converges to the unique optimal solution of (1) closest to $x_{0}$ as $\eta\to\infty$ (Theorem 2.7). We also give tight bounds on the objective value of the solution obtained using finite $\eta$ (Lemma 2.4).
Beyond linear objectives, we investigate the use of projected subgradient ascent for maximizing convex functions. In the convex minimization setting, projected subgradient descent requires vanishing step sizes. In contrast, we show that for convex maximization, projected subgradient ascent converges to a first-order stationary point when using arbitrarily large step sizes (Theorem 3.14). This result holds for continuous convex functions in infinite-dimensional Hilbert spaces. Previous work on nonconvex minimization in finite-dimensional spaces (e.g., [2]) typically rely on differentiability, Lipschitz continuity of the gradient, and the Kurdyka-Łojasiewicz inequality to ensure convergence. Our analysis removes these structural requirements by leveraging convexity of the function being maximized.
In practice, the large step size regime of projected subgradient ascent may lead to faster convergence and yield meaningful behavior. In the limit as the step size goes to infinity, projected subgradient ascent leads to the conditional gradient method with unit step size (Section 3.3). This limit also generalizes the iterated linear optimization paradigm introduced in [13].
The idea of using a single projection for linear optimization was previously considered in [15] in the context of linear programming where $S$ is a polytope. The same procedure was rediscovered in [19] under the assumption of strict complementarity, and subsequently extended to infinite-dimensional convex problems satisfying a sharpness property in [9]. Here we consider general convex sets that may have smooth boundaries. Such sets arise in a variety of settings including in semidefinite programming. The idea of using a single projection for linear optimization was also considered recently, and independently of our work, in [22]. Compared to [22], the bounds we derive in Section 2 are tighter, and we prove convergence to a unique optimum solution besides proving convergence of the objective value as was done in [22].
@@
[10] undertook a complexity analysis and show that on many domains—the simplex, $\ell_{p}$-balls, nuclear-norm ball, the flow polytope, Birkhoff polytope, and permutahedron—the best known algorithms for linear optimization are asymptotically faster than the best known algorithms for orthogonal projection. The reduction of linear optimization to orthogonal projection gives further evidence that linear optimization over a convex set is no harder than projecting to the same set. While [10] and [22] suggest this as a negative result for projection-based methods, we emphasize that the reduction can be used to obtain positive results, in that efficient algorithms for projections can lead to efficient algorithms for linear optimization.
@@ 2. Single Projection for Linear Optimization
Since $S$ is closed and convex, and $\langle c,x\rangle$ is a linear function of $x$, the set of maximizers $\mathcal{M}(c)$ is non-empty, closed, and convex. Let $x_{0}\in\mathcal{H}$. Denote by $\|\cdot\|$ the norm induced by the inner product. By the Hilbert Projection Theorem, there is a unique element $x^{*}\in\mathcal{M}(c)$ closest to $x_{0}$,
@@
[…]
@@
The use of a single projection for linear optimization was considered in [15] and [19] in the context of linear programming, where projecting a suitably rescaled cost vector yields an optimal basic feasible solution. [9] extended these results to infinite-dimensional spaces under a sharpness assumption. While these work sought an explicit $\eta$ that produces an exact solution in the polyhedral setting, we consider more general (non-polyhedral) closed convex sets in a Hilbert space and derive bounds on $\eta$ that ensure an arbitrarily close approximation to the objective value. Note that for smooth sets, no finite $\eta$ achieves optimality.
A result similar to Lemma 2.4 appears in [22]. However, our lemma provides a more precise characterization that depends on $x_{0}$. This also allows us to prove convergence of $P_{S}(x_{0}+\eta c)$ in Theorem 2.7, including in the case of general (non-polyhedral) convex sets $S$.
@@
The next key lemma shows $\langle c,x^{\eta}\rangle$ approaches $\langle c,x^{*}\rangle$ as $\eta$ grows.
@@
Indeed, let $c=(\tfrac{1}{2\eta},1)$. The optimal solution is $x^{*}=(1,1)$ while $x^{\eta}=(\tfrac{1}{2},1)$.
Figure 2. Illustration of Example 2.6. By selecting $\eta$ sufficiently large we can ensure $\langle c,x^{*}\rangle-\langle c,x^{\eta}\rangle\leq\epsilon$ independent of $c$, while $\|x^{*}-x^{\eta}\|$ remains large.
@@
for all $\eta>0$. Therefore, $x^{\eta}$ lies in a closed ball of radius $\|x^{*}\|$ centered at $x_{0}$. Closed balls in Hilbert space are weakly sequentially compact. Consider an arbitrary increasing sequence $\{\eta_{j}\}_{j\in\mathbb{N}}$ such that $x^{\eta_{j}}\rightharpoonup\overline{x}$ for some $\overline{x}\in S$. We show that $x^{\eta_{j}}\to x^{*}$ as $j\to\infty$.
@@
Lemma 2.4 implies $\|x^{\eta}-x_{0}\|^{2}\leq\|x^{*}-x_{0}\|^{2}$. Therefore $\limsup_{j}\|x^{\eta_{j}}-x_{0}\|^{2}\leq\|x^{*}-x_{0}\|^{2}$. The weak lower semicontinuity of the norm yields $\|x^{*}-x_{0}\|\leq\liminf_{j}\|x^{\eta_{j}}-x_{0}\|$. Since the limit inferior is at most the limit superior, combining both inequalities, we obtain
@@ 3. Projected Subgradient Ascent
In this section, we analyze the behavior of projected subgradient ascent for maximizing a continuous convex function $f$ over a closed convex set $S\subseteq\mathcal{H}$. The problem of maximizing a convex function is NP‐hard (e.g., [20]). Therefore, we focus on obtaining first-order stationary points.
@@
converges to a global minimizer (see, e.g., [7]). In contrast, when maximizing a convex function using projected subgradient ascent (PGA) convergence to a global optimum is not guaranteed, regardless of the choice of step sizes. Rather than vanishing steps, we focus on the case of arbitrarily large step sizes and show that PGA always converges to a first-order stationary point of $f$.
@@
Proposition 3.9 shows that in the convex maximization regime with PGA, the sequence $\{f(x_{k})\}$ is always nondecreasing. When $f$ is bounded, this implies $\{f(x_{k})\}$ converges. If the step sizes are bounded from above, Theorem 3.12 shows that the accumulation set of $\{x_{k}\}$ is connected. Moreover, if the step sizes are bounded from below, Theorem 3.14 shows all of the accumulation points of $\{x_{k}\}$ are first-order stationary points of $f$.
These results illustrate a contrast between convex minimization with PGD and convex maximization with PGA. The convex minimization setting requires vanishing steps for convergence, while in the convex maximization setting, convergence is guaranteed with large steps. In Section 3.3, we consider the limiting behavior when all of the step sizes go to infinity.
@@ 3.1. Linear Functions
In the general case of possibly infinite-dimensional spaces, weak convergence under arbitrary step sizes can be shown using notions from monotone operator theory (see, e.g., [5]). We formally state this known result (see, e.g., Theorem 2.5 in [4]) as Theorem 3.7 to build the intuition for taking large step sizes before we tackle the broader class of general convex functions.
@@
Let $\{\eta_{k}\}_{k\in\mathbb{N}}$ be a sequence of step sizes such that $\sum_{k=0}^{\infty}\eta_{k}=\infty$. Let $\{x_{k}\}_{k\in\mathbb{N}}$ be a sequence of iterates generated by PGA with $f(x)=\langle c,x\rangle$. The sequence $\{x_{k}\}_{k\in\mathbb{N}}$ converges weakly to a point in $\mathcal{M}(c)$.
@@
We emphasize the difference between the results in Theorem 2.7 and Theorem 3.7. The first theorem analyzes the behavior of a single step of projected gradient ascent $P_{S}(x_{0}+\eta c)$ as the step size $\eta$ goes to infinity, proving strong convergence to the unique solution in $\mathcal{M}(c)$ closest to $x_{0}$. In contrast, the second theorem establishes weak convergence of projected gradient ascent to some solution in $\mathcal{M}(c)$ after an infinite number of finite steps.
@@ 3.2. Convex Functions
Now, we consider the case of maximizing a continuous convex function with PGA.
@@
Previous results that apply to the setting of maximizing convex functions require significantly stronger assumptions. For example, the convergence results in [2] apply in finite-dimensional spaces when $f$ is differentiable with a Lipschitz continuous gradient over a nonempty closed set $S$, and satisfies the Kurdyka-Łojasiewicz property.
@@
(2) $f_{k}(x_{k+1})\geq f_{k}(x_{k}).$
Note that $x_{k+1}=P_{S}(x_{k}+\eta_{k}g_{k})$ and $x_{k}=P_{S}(x_{k})$. By Proposition 2.8 with $\eta_{1}=0$ and $\eta_{2}=\eta_{k}$ we know that $\langle g_{k},x_{k+1}\rangle\geq\langle g_{k},x_{k}\rangle$. This implies $\langle x_{k+1}-x_{k},g_{k}\rangle\geq 0$ and now (2) follows from the definition of $f_{k}$.
@@
Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\to\mathbb{R}$ be continuous and convex, $S\subseteq\mathcal{H}$ be a compact and convex set, and $\{x_{k}\}_{\mathbb{N}}$ be a sequence generated by PGA with step sizes $\{\eta_{k}\}_{\mathbb{N}}$ where $\limsup_{k\to\infty}\eta_{k}<\infty$. Then the set of accumulation points of $\{x_{k}\}_{\mathbb{N}}$ is connected.
@@
Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\to\mathbb{R}$ be continuous and convex, $S\subseteq\mathcal{H}$ be a compact and convex set, and $\{x_{k}\}_{\mathbb{N}}$ be a sequence generated by PGA with step sizes $\{\eta_{k}\}_{\mathbb{N}}$ where $\limsup_{k\to\infty}\eta_{k}<\infty$ and $\liminf_{k\to\infty}\eta_{k}>0$. Then every accumulation point of $\{x_{k}\}_{\mathbb{N}}$ is a first-order stationary point of $f$.
@@ 3.3. Conditional Gradient and Iterated Linear Optimization
$x_{k+1}\in\operatorname*{\arg\!\max}_{z\in S}\,\langle x_{k},z\rangle,$
@@
Acknowledgements
This research was partially funded by the Brown University Advanced Undergraduate Research Fellowship under the SPRINT/UTRA program.